OpenStudy (anonymous):

A window washer pulls herself upward using the bucket-pulley apparatus. The mass of the person plus the bucket is 65 kg. (a) How hard must she pull downward to raise herself slowly at constant speed? (b) If she increases this force by 10%, what will her acceleration be?

OpenStudy (anonymous):

Consider the free body diagrams as shown in the file attached. Let the acc of system is a. For the washer T+N-m1g=m1a ............................(1) For the bucket T-N-m2g=m2a ............................(2) (a) a=0 Eqns (1) and (2) become T+N-m1g=0 T-N-m2g=0 Adding and rearranging, T=(m1+m2)/2g T=325N (b) We are specified the value of T here as 325+32.5=357.5N eqns (1)and (2) become 357.5+N-m1g=m1a 357.5-N-m2g=m2a Adding, 715-(m1+m2)g=(m1+m2)a Solve for 'a'

OpenStudy (anonymous):

thanks :)